OpenStudy (anonymous):

I need to know how to do this problem: Joan wants to jog 13 miles on a circular track 1/8 miles in diameter. (use 3.14 for pi)

OpenStudy (anonymous):

The question is?

OpenStudy (anonymous):

I'll try it again how many times must she circle the track

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

yeah where does the 12 come from

OpenStudy (anonymous):

= 13/ (Pi/8) ~ 33.1 circles

OpenStudy (anonymous):

It's just typo!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I think this math stuff is going to be the death of me

OpenStudy (anonymous):

Since the perimeter of circle is D * pi = pi/8

OpenStudy (anonymous):

Don't worry, you have entire OS building behind you :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

So many person thought as you do, but just within 2 weeks they send thanks you around showing how they ace math!

OpenStudy (anonymous):

The more you post, the more you learn :)

OpenStudy (anonymous):

I still don't think I understand how this problem is done what are you multiplying

OpenStudy (anonymous):

Suppose you can straight out the perimeter into a straight line! since Joan want to run total 13 miles around the circular track, => divide total with the perimeter. P = D pi = pi/8 miles

OpenStudy (anonymous):

DIVIDE, not mutiply! => Total length/ P = 13 / (pi/8) = 33 rounds.